OpenStudy (anonymous):

PLEASE HELP!!!!! a 0.0500 g mixture of Cu2O and CuO contains 0.425 g Cu. What is the mass of CuO in the mixture?

OpenStudy (aaronq):

you have to use two equations: (1) mass of CuO(molar mass of Cu) + mass of Cu2O (mass of 2 Cu) = 0.425 g (2) mass of CuO (molar mass of CuO) + mass of Cu2O (molar mass of Cu2O)= 0.5g you can use x and y instead of words (1) x(molar mass of Cu) + y(mass of 2 Cu) = 0.425 g (2) x(molar mass of CuO) + y(molar mass of Cu2O)= 0.5g x is the mass of CuO

OpenStudy (aaronq):

in the first equation i meant (1) x(molar mass of Cu) + y(MOLAR** mass of 2 Cu) = 0.425 g